OpenStudy (anonymous):

Find the width of a rectangle whose diagonal is 25 ft and whose length is 12 ft

OpenStudy (anonymous):

X^2 + y^2 =C^2 12^2 + x^2 = 25^2 144+ x^2 = 625 x^2= 481 x= sqrt(481)